# REACT3D: Recovering Articulations for Interactive Physical 3D Scenes

### v1

@@ IV-A Experimental Setup
Datasets
To comprehensively evaluate our method and baselines, we use ScanNet++ [2], a widely adopted dataset that provides high-quality posed RGB-D reconstructions across diverse indoor environments. We select 30 scenes rich in objects with potential articulation, ranging from single rooms to full apartments. For quantitative evaluation, we adopt Articulate3D [12] as ground truth, which augments ScanNet++ with part-level semantic labels and articulation annotations (joint type, origin, axis, and motion range). Our evaluation focuses on objects with open-close motion, while small controls such as buttons and knobs are excluded. For benchmarking, we evaluate our method against the baselines URDFormer and DRAWER.
Evaluation metrics
We assess performance along two dimensions: movable-part detection and articulation parameters. For detection, we compute vertex-level IoU over all pairs of predicted movable-part meshes and ground-truth part meshes from Articulate3D in world coordinates. As the predicted parts are segmented directly from the input static scene, both predictions and ground truth reside in the same coordinate frame, requiring no additional alignment. A prediction is counted as a true positive if its IoU with at least one ground-truth part exceeds $\tau$. We report precision, recall, and F1 at $\tau\in\{0.25,0.50\}$ across all scene objects.
@@
Input
DRAWER (GT-mesh)
REACT3D
Figure 6: Qualitative comparisons for scene-level movable-part detection and articulation estimation. Red arrows indicate revolute joints, while blue arrows indicate prismatic joints. In (a), our method achieves significantly higher performance than DRAWER when openable objects lack visible handles. In (b), DRAWER misclassifies drawers as revolute joints, whereas our method provides correct predictions. Visualizations are generated with Open3D [34].
@@
Method
#Openable
Objects
$\tau^{*}=0.25$ $\tau^{*}=0.50$
@@
URDFormer 226 $50.9$ $0.583$ 0.500 $22.6$ $11.5$ $11.5$ $1.3$ $46.5$ $0.491$ 0.609 $19.0$ $8.9$ $8.9$ $1.8$
DRAWER (GT-mesh) 540 $73.7$ $0.288$ $1.621$ $14.1$ $10.4$ $10.0$ $5.2$ $78.9$ $0.313$ $1.714$ $10.6$ $8.3$ $8.0$ $3.7$
Ours 540 91.5 0.203 $1.161$ 39.3 35.9 35.9 24.8 92.2 0.180 $1.130$ 35.7 33.0 33.0 23.5
@@ IV-B Baselines
URDFormer
URDFormer generates an interactive scene from a single RGB image without guaranteed metric scale or an explicit mapping to the source 3D scene. For fair comparison, we adopt a concise placement-scaling protocol. For each static scene, we select a region rich in potentially interactive objects and capture a posed RGB-D image. We run the official model, extract per-object 3D bounding boxes and a global box, estimate a global scale $s$ as the ratio between the vertical extent of the ground-truth region and that of the predicted global box, then scale and place the prediction back into the static scene using the pose and depth.
As the output is not segmented from the scene mesh, vertex-level IoU is inapplicable. We therefore evaluate detection via 3D coverage: for a predicted box $B$ and a ground-truth movable-part mesh $M$ with vertices $V(M)$ in world coordinates, $\mathrm{cov}(B,M)=\frac{|V(M)\cap B|}{|V(M)|}$ and a part is detected if $\max_{B}\mathrm{cov}(B,M)\geq\tau$. Articulation is then evaluated with the same protocol and metrics defined above.
DRAWER
DRAWER reconstructs a scene mesh from RGB frames via monocular depth and normal estimation and generates an interactive scene on top. For fair comparison on ScanNet++, we bypass reconstruction and directly use ground-truth meshes as input. All downstream modules follow official settings and pretrained weights, except that the original GPT-4o component is replaced with the same LLaVA model used in our framework. We denote this configuration as DRAWER (GT-mesh). This protocol aligns the input domain with ours, ensuring independence from reconstruction errors in evaluation.
@@ IV-C Results
Openable object detection
We first assess how many openable objects each method recovers in the generated interactive scenes. Table I compares URDFormer, DRAWER (GT-mesh), and REACT3D under the experimental setup. Note that the output scenes of URDFormer are not at true scale as explained in Sec. IV-B, whereas the other methods produce scenes at true scale. Across all metrics, our framework achieves the best results and substantially outperforms both URDFormer and DRAWER, demonstrating its superior ability to consistently and accurately recover interactive objects across diverse indoor scenes.
Articulation estimation
As shown in Tab. II, our approach achieves over 90% joint-type accuracy, substantially surpassing all baselines. It also attains the lowest MD and OE among methods with learned articulation, outperforming DRAWER, the prior state of the art with true-scale results. Although URDFormer reports the lowest OE, this is due to its heuristic strategy of presetting articulation axes to vertical or horizontal relative to the ground based on joint type. Fig. 6 shows scene-level comparisons of movable-part detection and articulation results, while Fig. 7 compares interactive scenes generated by the native methods.
We further report results under the MOD setting in Tab. II, following the same experimental setup. This setting provides a cumulative view of articulation parameter correctness and reveals which metric constitutes the primary performance bottleneck. Here, PDet measures detection accuracy, while +M, +MO, and +MOD denote progressively stricter criteria: correct motion type, motion type with OE $10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, and additionally MD $0.25\text{\,}\mathrm{m}$\$, respectively [7]. Each stage is evaluated only if all preceding conditions are satisfied.
Ablation study
To assess the contribution of our articulation refinement module, we conduct an ablation study by comparing results with and without it. As shown in Tab. III, refinement consistently improves performance across both IoU thresholds. The reduction in minimum distance demonstrates more accurate joint localization, while the dramatic decrease in orientation error underscores its importance for reliable axis direction estimation. Overall, the results demonstrate that refinement is crucial for producing precise articulation parameters and achieving robust interactive scene generation.
TABLE III: Ablation study on articulation refinement. We report minimum distance (MD) and orientation error (OE) at IoU thresholds $\tau\!\in\!\{0.25,0.50\}$. Refinement consistently enhances accuracy, with notably large improvements in OE.
@@
w/o refinement $0.226$ $6.721$ $0.213$ $6.638$
w/ refinement 0.203 1.160 0.180 1.130
@@ V Conclusion
We presented REACT3D, a scalable zero-shot framework that transforms static 3D scenes into interactive digital twins with articulated objects. By combining open-vocabulary detection, articulation refinement, hidden-geometry completion, and seamless simulator integration, REACT3D achieves state-of-the-art performance on scene-level openable-object detection and articulation estimation. Our results demonstrate the practical utility of REACT3D in generating simulation-ready assets, providing a foundation for large-scale research in articulated scene understanding and embodied intelligence.

### v3

@@ IV-A Experimental Setup
Datasets.  To comprehensively evaluate our method and baselines, we use ScanNet++ [31], a widely adopted dataset that provides high-quality posed RGB-D reconstructions across diverse indoor environments. We select 30 scenes rich in articulated objects, ranging from single rooms to full apartments. For quantitative evaluation, we adopt Articulate3D [8] as ground truth, which augments ScanNet++ with part-level semantic labels and articulation annotations (joint type, origin, axis, and motion range). Our evaluation focuses on objects with open-close motion, while small controls such as buttons and knobs are excluded. For benchmarking, we evaluate our method against the baselines URDFormer and DRAWER.
Evaluation metrics.  We assess performance along two dimensions: movable-part detection and articulation parameters. For detection, we compute vertex-level IoU over all pairs of predicted movable-part meshes and ground-truth part meshes from Articulate3D in world coordinates. As the predicted parts are segmented directly from the input static scene, both predictions and ground truth reside in the same coordinate frame, requiring no additional alignment. A prediction is counted as a true positive if its IoU with at least one ground-truth part exceeds $\tau$. We report precision, recall, and F1 at $\tau\in\{0.25,0.50\}$ across all scene objects.
@@
Dataset/Method #Openable Objects $\tau^{*}=0.25$ $\tau^{*}=0.50$
@@
URDFormer 226 $51.0$ $0.583$ 0.50 $22.6$ $11.5$ $11.5$ $2.2$ $46.5$ $0.491$ 0.61 $19.0$ $8.8$ $8.8$ $3.1$
DRAWER (GT-mesh) 540 $73.7$ $0.288$ $1.62$ $14.1$ $10.4$ $10.0$ $5.2$ $78.9$ $0.313$ $1.71$ $10.6$ $8.3$ $8.0$ $3.7$
REACT3D 540 91.5 0.203 $1.16$ 39.3 35.9 35.9 25.2 92.2 0.180 $1.13$ 35.7 33.0 33.0 23.9
(a) Static Scan
(b) Output
(c) Revolute
(d) Prismatic
Figure 8: Real-world Experiment. Interactive scene generated from a real-world scan, enabling robotic manipulation of revolute and prismatic joints.
Implementation Details.  The threshold values utilized in our pipeline were determined empirically and stay static across all experiments. We conducted a sensitivity analysis on the parameters of the non-learning algorithms, which demonstrates the robustness of our parameter choices.
@@ IV-B Baselines
URDFormer [1] generates an interactive scene from a single RGB image without guaranteed metric scale or an explicit mapping to the source 3D scene. For fair comparison, we adopt a concise placement-scaling protocol. For each static scene, we select a region rich in potentially interactive objects and capture a posed RGB-D image. We run the official model, extract per-object 3D bounding boxes and a global box, estimate a global scale $s$ as the ratio between the vertical extent of the ground-truth region and that of the predicted global box, then scale and place the prediction back into the static scene using the pose and depth. As the output is not segmented from the scene mesh, vertex-level IoU is inapplicable. We therefore evaluate detection via 3D coverage: for a predicted box $B$ and a ground-truth movable-part mesh $M$ with vertices $V(M)$ in world coordinates, $\mathrm{cov}(B,M)=\frac{|V(M)\cap B|}{|V(M)|}$ and a part is detected if $\max_{B}\mathrm{cov}(B,M)\geq\tau$. Articulation is then evaluated with the same protocol and metrics defined above.
DRAWER [30] reconstructs a scene mesh from RGB frames via monocular depth and normal estimation and generates an interactive scene on top. For fair comparison on ScanNet++, we bypass reconstruction and directly use ground-truth meshes as input. All downstream modules follow official settings and pretrained weights, except that the original GPT-4o component is replaced with Gemini 2.5 Flash. We denote this configuration as DRAWER (GT-mesh). This protocol aligns the input domain with ours, ensuring independence from reconstruction errors in evaluation.
@@ IV-C Results
Openable object detection.  We first assess how many openable objects each method recovers in the generated interactive scenes. Table I compares URDFormer, DRAWER (GT-mesh), and REACT3D under the experimental setup. Note that the output scenes of URDFormer are not at true scale as explained in Sec. IV-B, whereas the other methods produce scenes at true scale. Across all metrics, our framework achieves the best results and substantially outperforms both baselines, demonstrating its superior ability to consistently and accurately recover interactive objects across diverse indoor scenes.
Articulation estimation.  As shown in Tab. II, our approach achieves over 90% joint-type accuracy, substantially surpassing all baselines. It also attains the lowest MD and OE among methods with learned articulation, outperforming DRAWER, the prior state of the art with true-scale results. Although URDFormer reports the lowest OE, this is due to its heuristic strategy of presetting articulation axes to vertical or horizontal relative to the ground based on joint type. Fig. 3 shows scene-level comparisons of movable-part detection and articulation results, while Fig. 7 compares interactive scenes generated by native methods.
We further report results under the MOD setting in Tab. II, following the same experimental setup. This setting provides a cumulative view of articulation parameter correctness and reveals which metric is the primary performance bottleneck. Here, PDet measures detection accuracy, while +M, +MO, and +MOD denote progressively stricter criteria: correct motion type, motion type with OE $10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, and additionally MD $0.25\times$ the part mesh diagonal, respectively [18]. Each stage is evaluated only if all preceding conditions are satisfied.
Ablation study.  To assess the contribution of our articulation refinement module, we conduct an ablation study by comparing results with and without it. As shown in Tab. III, refinement consistently improves performance across both IoU thresholds. The reduction in minimum distance demonstrates more accurate joint localization, while the dramatic decrease in orientation error underscores its importance for reliable axis direction estimation. Overall, the results demonstrate that refinement is crucial for producing precise articulation parameters and achieving robust interactive scene generation.
TABLE III: Ablation study on articulation refinement. minimum distance (MD) and orientation error (OE) are reported at $\tau\!\in\!\{0.25,0.50\}$. Refinement consistently enhances accuracy, with notably large improvements in OE.
@@
w/o refinement $0.226$ $6.72$ $0.213$ $6.64$
w/ refinement 0.203 1.16 0.180 1.13
Runtime analysis.  Our framework demonstrates significantly higher computational efficiency compared to the official DRAWER baseline. In a representative scene, our approach requires 33,500s to process the full RGB-D sequence, whereas DRAWER takes 43,638s, which is $\sim$30% longer. Notably, our current sequential implementation offers substantial speedup potential through parallelized modular processing. Further analysis reveals that runtime is highly correlated with the total frame count. Consequently, implementing keyframe extraction can significantly reduce the runtime without compromising the quality of the results. This is evidenced by our real-world application scenario (Fig. 8), where using only keyframes achieved a runtime of 3,411s, significantly lower than processing full sequences from standard datasets.
Real-world validation.  We further evaluate our pipeline by converting real-world scans into interactive digital twins for robotic manipulation. As shown in Fig. 8, a mobile manipulator successfully interacts with diverse articulated objects leveraging the recovered parameters. This demonstrates that our interactive representations are sufficiently accurate and physically grounded, enabling seamless integration with downstream robotic applications.
@@ V Conclusion
We present REACT3D, a scalable zero-shot framework that transforms static 3D scenes into interactive digital twins with articulated objects. By combining open-vocabulary detection, articulation refinement, hidden-geometry completion, and seamless simulator integration, REACT3D achieves state-of-the-art performance on scene-level openable-object detection and articulation estimation. Our results demonstrate the practical utility of REACT3D in generating simulation-ready assets, providing a foundation for large-scale research in articulated scene understanding and embodied intelligence.